OpenStudy (i<3mybaby):

Use the given formula to solve the following problem: When the amount financed is $15,480, the monthly payment is $621.33, and the monthly interest rate is 0.00583, how long do you have to pay? A. 18 months B. 23 months C. 27 months D. 32 months @dan815

OpenStudy (valpey):

Are you using a financial calculator, or are you doing this by hand?

OpenStudy (i<3mybaby):

By hand

OpenStudy (dan815):

use a calculator

OpenStudy (i<3mybaby):

I tried that its not working

OpenStudy (i<3mybaby):

I cant figure it out :(

OpenStudy (dan815):

ok hang on

OpenStudy (i<3mybaby):

K

OpenStudy (i<3mybaby):

Arm thanks but that helps me out none at this point cause I dont knownthe order

OpenStudy (aum):

P = $621.33 A = $15,480 r = 0.00583 Plug these three numbers into the formula.

OpenStudy (i<3mybaby):

Dan just pm me

OpenStudy (i<3mybaby):

@dan815 ^^^

OpenStudy (i<3mybaby):

I figured it out

OpenStudy (dan815):

This is the equation i found its kinda long, bear with me P=15480 a=621.33 r=0.00583

OpenStudy (i<3mybaby):

Its ok Ps I still haven't event got to the q

OpenStudy (i<3mybaby):

So hence y I said pm me @dan815

OpenStudy (dan815):

ok sry internet crashed

OpenStudy (dan815):

P=Payment to be made in total A=Monthly installements for the payment R=rate of interset per month, that is adde to the money that is left over to pay

OpenStudy (dan815):

1st month before the interest rate kicks in she will pay A, so P-A left over to pay P-A 2 nd month, now there is an interest on the payment left over and then A is subtracted (P-A)+r(P-A) - A, =(P-A)*(1+r) -A =y*k +y-p, if P-A=y , 1+r =k 3rd Month =(yk+y-p)*k +y-p =k^2y+yk+y-kp-p 4th month =k^3y+k^2y+ky+y-k^2p=kp-p ... n+1 months =k^ny+k^n-1y....k^0y-k^(n-1)p-..-k^0p =y((k^(n+1)-1)/(k-1))-p((k^n-1)/(k-1) =0 <--- equal to zero before after n months we have 0 payment left